# Robust Utility Optimization via a GAN Approach

### v1

@@ A.2 Additive and Multiplicative Penalization for the Volatility
We define the additive penalty function appearing in Equations 5 and 14
@@
An analytic solution of the robust utility optimization problem with instantaneous penalties building on $F_{m}$ can be derived along the same lines as in Section 4.1, by solving the equation system
@@
A.3 Settings with Analytic Solution
@@ A.3.1 Additional Results: Stock Market with 2 Uncorrelated Risky Assets – Robust Volatility Case
In this section, we analyze additional metrics for the settings (S) and (AS) of Section 4.1.2.
@@
We observed in Section 4.1.2, that for additive penalties involving $F_{a}$, the learned strategies $\pi^{\text{NN}}$ achieved $\operatorname{err}_{rel}^{\Sigma^{\star},\mu}(\pi^{\text{NN}},\pi^{\star})<0.2\%$ in both settings (S) and (AS). Moreover, the GAN training solves the optimization problem (15) similarly well also for the multiplicative penalization involving $F_{m}$ with $\operatorname{err}_{rel}^{\Sigma^{\star},\mu}(\pi^{\text{NN}},\pi^{\star})<0.2\%$ in all tested hyper-parameter combinations.
@@
We test two noise levels $\epsilon\in\{0.025,0.15\}$. For both (S) and (AS), both noise levels and both penalization functions $F_{a}$ and $F_{m}$, average expected utilities based on noisy market volatilities $\sigma^{noisy}$ for learned strategies and for analytic strategies are indistinguishable (on a 5% level of confidence).
@@
The oracle solution leads to an average expected utility that is significantly (on a 5% confidence level) higher than the one of either the learned or the analytic strategies. In particular, this additional information about the market leads to a significant advantage that manifests in larger expected utilities. This observation reinforces our choice, not to use the discriminator output as input for the generator (and vice versa) as it was done in Guo et al. (2022), since this should lead to unrealistically high expected utility results that can not be reproduced in a real world setting.
@@ A.3.2 Additional Results: Stock Market with 2 Correlated Risky Assets – Robust Volatility Case
We proceed to analyze additional metrics for the settings (PS), (PAS), (NAS) of Section 4.1.2.
@@
We recall that we had small relative errors with the additive regularization functional $F_{a}$ in Section 4.1.2. With the multiplicative regularization functional $F_{m}$, training the GAN objective (15) seems harder (or further hyper parameter tuning might be needed): while in (PS), $E^{*}(\pi^{\text{NN}})$ is lower than $E^{*}(\pi^{*})$ on the third decimal place, for (PAS) and (NAS), their difference even ranges within first and second decimal places.
@@
A similar distinction in performance crystallizes for the noisy market metric, where we use the same noise levels as in Section A.3.1. When considering the additive regularization $F_{a}$, across all settings and for both noise levels, $\mathbb{E}_{\Sigma^{noisy}}\ \left[E_{u}(\pi^{\text{NN}},\Sigma^{noisy})\right]$ for learned strategies is indistinguishable (on a 5% level of confidence) from $\mathbb{E}_{\Sigma^{noisy}}\left[E_{u}(\pi^{*},\Sigma^{noisy})\right]$. Strategies trained with $F_{m}$ perform significantly worse in $E_{u}$ than the analytic strategies based on noisy volatilities $\Sigma^{noisy}$. For (PS) and (PAS), this effect seems to weaken with higher noise. In (NAS), even in high-noise scenarios, the robustly learned strategy is outperformed by the analytic strategies in the metric $E_{u}$.
As in the uncorrelated setting, we also observe that in low-noise settings, $E_{u}(\pi,\Sigma^{noisy})$ is highest for non-robustly trained strategies $\pi$, while in settings with large noise, non-robustly trained $\pi$s result in smallest expected utility $E_{u}(\pi,\Sigma^{noisy})$.
@@
Also for correlated assets, $\mathbb{E}_{\Sigma^{noisy}}\left[E_{u}(\pi^{\text{oracle}},\Sigma^{noisy})\right]$ is significantly (on a 5% confidence level) higher than for either the learned or the analytic strategies.
@@ A.3.3 Additional Results: Stock Market with 5 Uncorrelated Risky Assets – Robust Volatility Case
We analyze the setting with 5 uncorrelated risky assets of Section 4.1.2.
@@
When considering the additive regularization $F_{a}$, for both noise levels, we have that $\mathbb{E}_{\Sigma^{noisy}}\ \left[E_{u}(\pi,\Sigma^{noisy})\right]$ for learned strategies $\pi^{\text{NN}}$ is indistinguishable (on a 5% level of confidence) from the analytic strategy $\pi^{*}$. Moreover, the results of comparing different penalization factors $\lambda_{1}$ are in line with intuition: for lower noise, $\lambda_{1}=10$ is the best choice in terms of $E_{u}$, for higher noise, we observe that $\lambda_{1}=0.1$ leads to highest $E_{u}$. The results are similar for regularization $F_{m}$, although it is not so clearly visible which value for $\lambda_{1}$ is optimal in low noise settings.
@@
Also here, the oracle strategy is significantly (on a 5% confidence level) better than either the learned or the analytic strategies. (For low-noise scenarios the difference is smaller than for high-noise scenarios.)
@@ A.4 Additional Results: Stock Market with Proportional Transaction Costs – Robust Volatility Case
We evaluate the settings of Section 4.1.2 with transaction costs in the noisy market metric. During the GAN training, we use $E_{u}$ on a validation set as lookback criterion for model selection for $\pi^{\text{NN}}$. This means that we sample a (reasonably large) number of validation paths of the risky assets based on realizations of noisy volatilities and choose $\pi^{\text{NN}}$ from our training history that maximized a sample estimate of $E_{u}$ on these validation paths. As in our test criterion, we consider for these noisy validation paths both the noise levels $\epsilon\in\{0.025,0.15\}$. This means that at testing time, we are able to compare for each scenario two NN-based strategies: one in which we stop the training based on a noisy estimate of $E_{u}$ with the correct market noise, and one where market noise is over- or underestimated.
@@ A.4.1 Stock Market with 2 Risky Assets
Finally, as expected, we observe that the NN-strategies selected based on the correct market noise perform generally better than the ones selected based on under- or overestimated market noise. However, only in a minority of experiments we observe that $\pi^{\text{NN}}$ selected based on a wrong estimate of market noise performs worse than the analytic strategy.
@@ A.5 Additional Results: Robust Drift and other Utility Functions – No Transaction Costs
In general, whether or not robustifying the drift is beneficial in terms of $E_{u}$ depends on the choice of both $\lambda_{1},\lambda_{2}$, as we have already seen before in Section A.4.2.
For the market setting with (AS) and (AD) of Section 4.1.3 and for the case of power utility with $p=2$, we performed a grid search over $\lambda_{1},\lambda_{2}\in\{0.01,0.1,0.5,1,10,100\}$, with results shown in Figure 11, where we evaluated the average expected utility with noisy market drift and volatility. The best combination of scaling parameters yields $\mathbb{E}_{\Sigma^{noisy},\mu^{noisy}}\left[E_{u}(\pi,\Sigma^{noisy},\mu^{noisy})\right]\approx 0.15628$ For comparison, in the non-robust case and in the case where only $\Sigma$ was robustified (corresponding to $\lambda_{2}=\infty$), the (maximal) values of average $E_{u}$ were $0.01523$ and $0.01556$, respectively. In particular, robustifying both the volatility and drift is beneficial when choosing the appropriate penalty scaling factors for the given market situation. Moreover, we see in Figure 11 that the calibration of the scaling parameter for the drift is more crucial than for the volatility.

### v2

@@ A.2 Additive and Multiplicative Penalization for the Volatility
We define the additive penalty function appearing in Equations˜5 and 14
@@
An explicit solution of the robust utility optimization problem with instantaneous penalties building on $F_{m}$ can be derived along the same lines as in Section˜4.1, by solving the equation system
@@
A.3 Settings with Explicit Solution
@@ A.3.1 Additional Results: Stock Market with 2 Uncorrelated Risky Assets – Robust Volatility Case
In this section, we analyze additional metrics for the settings (S) and (AS) of Section˜4.1.2.
@@
We observed in Section˜4.1.2, that for additive penalties involving $F_{a}$, the learned strategies $\pi^{\text{NN}}$ achieved $\operatorname{err}_{rel}^{\Sigma^{\star},\mu}(\pi^{\text{NN}},\pi^{\star})<0.2\%$ in both settings (S) and (AS). Moreover, the GAN training solves the optimization problem (15) similarly well also for the multiplicative penalization involving $F_{m}$ with $\operatorname{err}_{rel}^{\Sigma^{\star},\mu}(\pi^{\text{NN}},\pi^{\star})<0.2\%$ in all tested hyper-parameter combinations.
@@
We test two noise levels $\epsilon\in\{0.025,0.15\}$. For both (S) and (AS), both noise levels and both penalization functions $F_{a}$ and $F_{m}$, average expected utilities based on noisy market volatilities $\sigma^{noisy}$ for learned strategies and for explicit strategies are indistinguishable (on a 5% level of confidence).
@@
The oracle solution leads to an average expected utility that is significantly (on a 5% confidence level) higher than the one of either the learned or the explicit strategies. In particular, this additional information about the market leads to a significant advantage that manifests in larger expected utilities. This observation reinforces our choice, not to use the discriminator output as input for the generator (and vice versa) as it was done in Guo et al. (2022), since this should lead to unrealistically high expected utility results that can not be reproduced in a real world setting.
@@ A.3.2 Additional Results: Stock Market with 2 Correlated Risky Assets – Robust Volatility Case
We proceed to analyze additional metrics for the settings (PS), (PAS), (NAS) of Section˜4.1.2.
@@
We recall that we had small relative errors with the additive regularization functional $F_{a}$ in Section˜4.1.2. With the multiplicative regularization functional $F_{m}$, training the GAN objective (15) seems harder (or further hyper parameter tuning might be needed): while in (PS), $E^{*}(\pi^{\text{NN}})$ is lower than $E^{*}(\pi^{*})$ on the third decimal place, for (PAS) and (NAS), their difference even ranges within first and second decimal places.
@@
A similar distinction in performance crystallizes for the noisy market metric, where we use the same noise levels as in Section˜A.3.1. When considering the additive regularization $F_{a}$, across all settings and for both noise levels, $\mathbb{E}_{\Sigma^{noisy}}\ \left[E_{u}(\pi^{\text{NN}},\Sigma^{noisy})\right]$ for learned strategies is indistinguishable (on a 5% level of confidence) from $\mathbb{E}_{\Sigma^{noisy}}\left[E_{u}(\pi^{*},\Sigma^{noisy})\right]$. Strategies trained with $F_{m}$ perform significantly worse in $E_{u}$ than the explicit strategies based on noisy volatilities $\Sigma^{noisy}$. For (PS) and (PAS), this effect seems to weaken with higher noise. In (NAS), even in high-noise scenarios, the robustly learned strategy is outperformed by the explicit strategies in the metric $E_{u}$.
[…]
@@
Also for correlated assets, $\mathbb{E}_{\Sigma^{noisy}}\left[E_{u}(\pi^{\text{oracle}},\Sigma^{noisy})\right]$ is significantly (on a 5% confidence level) higher than for either the learned or the explicit strategies.
@@ A.3.3 Additional Results: Stock Market with 5 Uncorrelated Risky Assets – Robust Volatility Case
We analyze the setting with 5 uncorrelated risky assets of Section˜4.1.2.
@@
When considering the additive regularization $F_{a}$, for both noise levels, we have that $\mathbb{E}_{\Sigma^{noisy}}\ \left[E_{u}(\pi,\Sigma^{noisy})\right]$ for learned strategies $\pi^{\text{NN}}$ is indistinguishable (on a 5% level of confidence) from the explicit strategy $\pi^{*}$. Moreover, the results of comparing different penalization factors $\lambda_{1}$ are in line with intuition: for lower noise, $\lambda_{1}=10$ is the best choice in terms of $E_{u}$, for higher noise, we observe that $\lambda_{1}=0.1$ leads to highest $E_{u}$. The results are similar for regularization $F_{m}$, although it is not so clearly visible which value for $\lambda_{1}$ is optimal in low noise settings.
@@
Also here, the oracle strategy is significantly (on a 5% confidence level) better than either the learned or the explicit strategies. (For low-noise scenarios the difference is smaller than for high-noise scenarios.)
@@ A.4 Additional Results: Stock Market with Proportional Transaction Costs – Robust Volatility Case
We evaluate the settings of Section˜4.1.2 with transaction costs in the noisy market metric. During the GAN training, we use $E_{u}$ on a validation set as lookback criterion for model selection for $\pi^{\text{NN}}$. This means that we sample a (reasonably large) number of validation paths of the risky assets based on realizations of noisy volatilities and choose $\pi^{\text{NN}}$ from our training history that maximized a sample estimate of $E_{u}$ on these validation paths. As in our test criterion, we consider for these noisy validation paths both the noise levels $\epsilon\in\{0.025,0.15\}$. This means that at testing time, we are able to compare for each scenario two NN-based strategies: one in which we stop the training based on a noisy estimate of $E_{u}$ with the correct market noise, and one where market noise is over- or underestimated.
@@ A.4.1 Stock Market with 2 Risky Assets
Finally, as expected, we observe that the NN-strategies selected based on the correct market noise perform generally better than the ones selected based on under- or overestimated market noise. However, only in a minority of experiments we observe that $\pi^{\text{NN}}$ selected based on a wrong estimate of market noise performs worse than the explicit strategy.
@@ A.5 Additional Results: Robust Drift and other Utility Functions – No Transaction Costs
In general, whether or not robustifying the drift is beneficial in terms of $E_{u}$ depends on the choice of both $\lambda_{1},\lambda_{2}$, as we have already seen before in Section˜A.4.2.
For the market setting with (AS) and (AD) of Section˜4.1.3 and for the case of power utility with $p=2$, we performed a grid search over $\lambda_{1},\lambda_{2}\in\{0.01,0.1,0.5,1,10,100\}$, with results shown in Figure 11, where we evaluated the average expected utility with noisy market drift and volatility. The best combination of scaling parameters yields $\mathbb{E}_{\Sigma^{noisy},\mu^{noisy}}\left[E_{u}(\pi,\Sigma^{noisy},\mu^{noisy})\right]\approx 0.15628$ For comparison, in the non-robust case and in the case where only $\Sigma$ was robustified (corresponding to $\lambda_{2}=\infty$), the (maximal) values of average $E_{u}$ were $0.01523$ and $0.01556$, respectively. In particular, robustifying both the volatility and drift is beneficial when choosing the appropriate penalty scaling factors for the given market situation. Moreover, we see in Figure˜11 that the calibration of the scaling parameter for the drift is more crucial than for the volatility.